make x the subject of the formula
5(x-3)=y(4-3x)

Answer by Alan3354(69443)

make x the subject of the formula
5(x-3)=y(4-3x)
5x-15 = 4y - 3xy
Collect x terms.
5x + 3xy = 4y + 15
x*(5 + 3y) = 4y + 15
x = (4y+15)/(3y+5)